Differentiate w.r.t x

y= √{x+√(x+√x)}
1
Expert's answer
2020-03-26T11:03:06-0400

"y=\sqrt{x+\sqrt{x+\sqrt{x}}}\\
\text{We apply chain rule while differentiating}\\
\frac{dy}{dx}=\frac{d}{dx}(\sqrt{x+\sqrt{x+\sqrt{x}}})\\
\text{We know that }\frac{d}{dx}(\sqrt{g(x)})=\frac{1}{2\sqrt{g(x)}} \frac{d}{dx}(g(x))\\
\text{So, we have}\\
 \frac{dy}{dx}=\frac{d}{dx}(\sqrt{x+\sqrt{x+\sqrt{x}}})\\
\;\;\;\;\;=\frac{1}{2\sqrt{x+\sqrt{x+\sqrt{x}}}}\frac{d}{dx}(x+\sqrt{x+\sqrt{x}})\\
\;\;\;\;\;=\frac{1}{2\sqrt{x+\sqrt{x+\sqrt{x}}}}(1+\frac{1}{2\sqrt{x+\sqrt{x}}}\frac{d}{dx}(x+\sqrt{x}))\\
\;\;\;\;\;=\frac{1}{2\sqrt{x+\sqrt{x+\sqrt{x}}}}(1+\frac{1}{2\sqrt{x+\sqrt{x}}}(1+\frac{1}{2\sqrt{x}}))\\"